THE ROTATIONAL SPECTROSCOPY OF FOUR ANTIMONY CONTAINING DIATOMIC MOLECULES: SbN, SbP, SbF and SbCl

\underline{S. A. COOKE}, and M. C. L. GERRY, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver, British Columbia, Canada, V6T 1Z1.

In separate experiments the molecules SbN, SbP, SbF and SbCl have been prepared by laser ablating antimony metal in the presence of NH_{3}, PH_{3}, SF_{6} or Cl_{2}, respectively. Their pure rotational spectra have been measured with a cavity pulsed jet Fourier transform microwave spectrometer. Hyperfine structure in the rotational spectra of these molecules has led to the determination of the ^{121}Sb and ^{123}Sb nuclear quadrupole coupling constants in all four molecules for the first time. This in turn has enabled the electronic structures of the Sb-pnictides and Sb-halides to be compared and contrasted. The measurement of the respective rotational constants has provided highly precise internuclear distances.